OpenStudy (anonymous):

What is the exact value of √36/2√6

OpenStudy (anonymous):

\[\sqrt{36}/2\sqrt{6}\]

OpenStudy (anonymous):

What's the square root of 36?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

So now your equation reads 6/2 sqrt6 right? Or are you supposed to divide 36 by 2 in the square root bar and then simplify?

OpenStudy (anonymous):

I've been on this test for 5 hours so I've lost focus. this is probably an easy question

OpenStudy (anonymous):

first one. its sqrt

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

its 6 over 2sqrt6

OpenStudy (anonymous):

3(sqrt6)=?

OpenStudy (anonymous):

no 2(sqrt6)

OpenStudy (anonymous):

its a radical expression not division

OpenStudy (anonymous):

√36/(2√6) = 6 / (2√6) = 3/√6